OpenStudy (anonymous):

verify the identity: 1+sec^2x/1+tan^2x=1+cos^2x

terenzreignz (terenzreignz):

This identity \[\Large \frac{1+\sec^2(x)}{1+\tan^2(x)}=1+\cos^2(x)\] ?

OpenStudy (anonymous):

its easy, just plug the values of sec^2 and tan^2 in terms of sin and cos and keep on simplifying, you'll get the value on the right hand side.

terenzreignz (terenzreignz):

Or use the fact that \[\Large 1+\tan^2(x) = \sec^2(x)\]

OpenStudy (anonymous):

good call!

OpenStudy (anonymous):

but that would be using another known identity? would it?

terenzreignz (terenzreignz):

We're supposed to know basic identities anyway. If you like, you can always start with the pythagorean identity and work from there. \[\Large \cos^2(x) +\sin^2(x) = 1\] and work from there. I avoid using the 'convert everything to sines and cosines' method to prove trigonometric identities if I can help it... it's just so... not elegant :D But effective. Still prefer turning one side into the other.... it's like MAGIC! :D